OpenStudy (anonymous):

Hello the average of three numbers is 10 solve using x+4x+4x+3

OpenStudy (anonymous):

Is this question even accurate?

OpenStudy (anonymous):

Yes it is. It says the average of three numbers is 10. The second is four times the first, and the third number is three more than the second.

OpenStudy (anonymous):

ok,then you probably wrote the expression wrong,when most people see this,they would probably immediately just simplify it and do it like an equation so you kind of have to rewrite it i think

OpenStudy (anonymous):

It is not

OpenStudy (anonymous):

The three numbers are 3, 12, and 15. Quite simply solved by adding the 3 numbers. x+3x+3x+3 = 9x+3 knowing the average is 10 means that (9x+3)/3 = 10 solve for x

OpenStudy (anonymous):

No reread the problem

OpenStudy (anonymous):

Oops

OpenStudy (anonymous):

I got the answer right, but I misprinted the relationship between the numbers. *x+3x+3x+3 should read x+4x+4x+3; this is indeed equal to 9x+3

OpenStudy (anonymous):

Yep I got 9x+3 too then I sub. 10 and that's where I got messed up. So its 9x+3/3=10 right?

OpenStudy (anonymous):

Thank you sir!

OpenStudy (anonymous):

Happy to help. Study hard!